The rate of return on Cherry Jalopies, Inc., stock over the last five years was 21 percent, 11 percent, -3 percent, 5 percent, and 8 percent. Over the same period, the return on Straw Construction Company’s stock was 16 percent, 22 percent, -2 percent, 5 percent, and 18 percent. Calculate the variances and the standard deviations for Cherry and Straw. ( Do not round intermediate calculations. Enter your variance as a decimal rounded to 5 decimal places. Enter your standard deviation as a percent rounded to 2 decimals.
Answer #1

Cherry Jalopies, Inc.:

Average Return = [0.21 + 0.11 + (-0.03) + 0.05 + 0.08] / 5
Average Return = 0.42 / 5
Average Return = 0.084 or 8.40%

Variance = [(0.21 - 0.084)^2 + (0.11 - 0.084)^2 + (-0.03 - 0.084)^2 + (0.05 - 0.084)^2 + (0.08 - 0.084)^2] / 4
Variance = 0.03072 / 4
Variance = 0.00768

Standard Deviation = (0.00768)^(1/2)
Standard Deviation = 0.0876 or 8.76%

Straw Construction Company:

Average Return = [0.16 + 0.22 + (-0.02) + 0.05 + 0.18] / 5
Average Return = 0.59 / 5
Average Return = 0.118 or 11.80%

Variance = [(0.16 - 0.118)^2 + (0.22 - 0.118)^2 + (-0.02 - 0.118)^2 + (0.05 - 0.118)^2 + (0.18 - 0.118)^2] / 4
Variance = 0.03968 / 4
Variance = 0.00992

Standard Deviation = (0.00992)^(1/2)
Standard Deviation = 0.0996 or 9.96%